What is the value of the finite population correction factor in the formula for o when

n=100 and N = 5000.


1
Expert's answer
2021-05-07T11:06:02-0400

Formula for finite population correction factor is given by,


"fpc = \sqrt{\frac{N-n}{N-1}}"


where,

sample size n = 100

population size N = 5000


Hence,

"fpc = \sqrt{\frac{5000-100}{5000-1}} = \sqrt\frac{4900}{4999} = 0.99004"